{Rm. NE43-236 \atop May 29, 1991}

John Hughes Attn: FPCA ’91 Dept. of Computing Science University of Glasgow Glasgow G12 8QQ Scotland

Dear Prof. Hughes,

I am enclosing the final copy of the paper “Incremental Polymorphism” (paper number 105) authored by myself and Prof. Nikhil for the FPCA ’91 conference proceedings. The paper has been type-set in Springer-Verlag’s LNCS format as instructed by you. Kindly acknowledge the receipt of the paper. My e-mail address is shail@abp.lcs.mit.edu.

Regards, Shail Aditya

encl: Paper – 27 pgs